# School of Information and Network Security

### 03 Mar 2024 v2

@@ 2.1 Large Language Model
To ensure that LLMs align with human preferences and values, Reinforcement Learning from Human Feedback (RLHF) strategies are often incorporated during the training process. RLHF employs reinforcement learning algorithms, such as Proximal Policy Optimization[], to adapt LLMs to human feedback. This method integrates human judgment into the training loop, aiming to develop LLMs that are better aligned with human values, exemplified by models like InstructGPT[4].
@@ 3 Method
In China, public safety constitutes the core of police work, encompassing critical areas such as crime prevention, case investigation, intelligence analysis, and the maintenance of public order. This significant responsibility is typically borne by various specialized police forces to ensure targeted and efficient maximization. For instance, narcs play an indispensable role in drug control efforts, with their primary tasks being the identification of individuals, locations, and items related to drug activities, thereby advancing drug control initiatives. Cyber police focus on maintaining order in cyberspace, analyzing and assessing the public’s emotional responses to social events, timely identifying and dealing with online rumors, and ensuring the stability and safety of the online environment. criminal police, security officers, and traffic police play crucial roles in the investigation of criminal cases, handling civil disputes, and the control of traffic violations, respectively. Furthermore, in response to the severe challenges posed by the emerging crime of telecommunications fraud, the criminal police department has established an anti-fraud center. Its main tasks include quickly filtering potential fraudulent information, effectively issuing warnings, and blocking telecommunications fraud activities to protect the public’s financial safety and reduce economic losses. This collaborative division of labor ensures the comprehensiveness and professionalism of public safety work, providing a solid foundation for maintaining social stability and citizen safety.
@@ 3.1 Tasks and Datasets
To comprehensively evaluate the performance of LLMs in the field of public security, CPSDBench has constructed a four-dimensional assessment task system encompassing text classification, information extraction, question answering, and text generation. These tasks incorporate a variety of datasets collected from actual public security business scenarios, thereby enabling a more thorough and in-depth assessment of LLMs’ understanding and generation capabilities concerning public security-related data. The evaluative content of CPSDBench and its correspondence to the police category in China are illustrated in the figure. Considering the high costs associated with utilizing commercial-grade LLMs (such as ChatGPT, ChatGLM-130B), the size of each dataset is typically limited to a range of several dozen to several hundred samples.
Figure 1: The evaluative content of CPSDBench and its correspondence to the police category in China.
@@ 3.1.2 Information Extraction
  • Drug-Related Case Reports Dataset: This dataset contains texts of reports related to drug cases, intended for named entity recognition and relationship extraction tasks. It includes five types of entities within the dataset. Notably, the data labels feature a significant number of duplicate entities, making the evaluation process more stringent and complex. A
    prediction is considered completely accurate only when it precisely extracts both the text and the quantity of entities.
@@ 3.1.3 Question Answering
  • Public Security Case Reading Comprehension Dataset: This includes objective and subjective reading comprehension questions. For objective comprehension, LLMs typically need to combine the material to provide a yes or no answer to the question. For subjective reading comprehension questions, LLMs are required to generate natural language text as the final answer.
@@ 3.1.4 Text Generation
  • Public Security Case Summary Dataset: This dataset focuses on creating concise and accurate summaries of public security cases, providing essential information in a structured and clear manner. It is designed to test the model’s ability to synthesize and summarize key case details efficiently.
@@ 3.2 Baselines and Prompt
Model Size (#P) URL Access mode
GPT-4 $\sim$ chat.openai.com API
ChatGLM 130B chatglm.cn API
MiniMax-ABAB-5.5 $\sim$ api.minimax.chat API
Baichuan2[29] $\sim$ www.baichuan-ai.com API
XVERSE 13B,65B chat.xverse.cn API
iFLYTEK Spark-3.0 $\sim$ xinghuo.xfyun.cn API
Qwen[30] 72B tongyi.aliyun.com/qianwen API
ERNIE Bot[31] $\sim$ yiyan.baidu.com API
Atom (LLaMA-2-Chinese) 1B,7B llama.family API
@@ 3.3 Evaluation Metrics
In the objective reading comprehension component of the question-answering task, this study adopts Accuracy as the primary evaluation metric to quantify the model’s performance in identifying the correct answers. For generative question-answering and text generation tasks, the study selects Bilingual Evaluation Understudy (BLEU) and Recall-Oriented Understudy for Gisting Evaluation (ROUGE) as assessment metrics. These metrics are utilized to evaluate the model’s performance in terms of the quality and accuracy of generated text. The specific computational processes are defined in Equations ( 7) and ( 8), respectively. Equation ( 7) elaborates on the calculation method for the BLEU metric, focusing primarily on the lexical overlap between the generated text and the reference text. Equation ( 8) describes the calculation process for the ROUGE metric, emphasizing the extent to which the generated text covers and retains the content of the original text. Together, these two metrics provide a comprehensive quantitative basis for evaluating the performance of models in generative question-answering and text generation tasks.Furthermore, we have also selected the BERT Score, calculated based on the BERT-base model, as the evaluation criterion for subjective reading comprehension questions. This approach aims to measure the semantic similarity between LLMs’ predictions and the labels from a semantic perspective.
@@ 4.1 Performance Evaluation
In this study, we conducted a comprehensive evaluation of advanced LLMs to explore their performance across different natural language processing tasks. The results are shown in Table 2-Table 5.Through systematic assessments in four key tasks: text classification, information extraction, question answering, and text generation, our aim was to unveil the strengths and limitations of mainstream LLMs in public security tasks.To achieve a more intuitive and comprehensive comparison of the performance of LLMs, this study also utilized radar charts to visually present the overall efficacy of the models, as shown in Figure 3.
@@
We further investigated the impact of language specificity on model performance. Notably, Chinese models such as ChatGLM-4, Minimax-abab, and Baichuan2 surpassed GPT-4 in text generation and question answering tasks. This result highlights the advantage of models pre-trained on extensive Chinese corpora in handling Chinese tasks, indicating that language-specific pre-training is crucial for enhancing the performance of LLMs in specific linguistic domains.
@@
Finally, input length is an important factor affecting the performance of LLMs in public safety tasks. We observed the variation in F1 scores of LLMs when processing texts of different lengths in the task of extracting information from drug-related cases, as illustrated in Figure 1. The analysis indicates that for the majority of models, the predictive capability of LLMs tends to decline with an increase in the length of the input text.
Figure 3: Overall performance comparison of LLMs. The text classification score is obtained by calculating the average accuracy across three subtasks. For information extraction, the F1 score from drug-related case tasks is used as the final score. The text generation task employs the BERT score as its final score.
@@
Model Sentiment Classification Rumor Detection Tel. Fraud Detection
GPT-4 0.7533 0.8975 0.8600
ChatGLM-4 0.7250 0.5750 0.8400
MiniMax-ABAB-5.5 0.7067 0.8375 0.8050
Baichuan2 0.6933 0.8350 0.8025
XVERSE-2-65B 0.7100 0.8175 0.4925
XVERSE-2-13B 0.6733 0.7475 0.6425
iFLYTEK Spark-3.0 0.6550 0.7175 0.0025
Qwen 0.4300 0.8500 0.7125
ERNIE Bot 0.5550 0.7575 0.6125
Atom-7B 0.4683 0.7050 0.6025
Atom-1B 0.0516 0.0600 0.4750
Table 3: The Results of LLMs in Information Extraction Task.
Model Precision Recall F1 Score
GPT-4 0.6721 0.6352 0.6295
ChatGLM-4 0.5939 0.5049 0.5136
MiniMax-ABAB-5.5 0.6108 0.5078 0.5159
Baichuan2 0.6566 0.5556 0.5644
XVERSE-2-65B 0.6778 0.5198 0.5526
XVERSE-2-13B 0.5937 0.4687 0.4942
iFLYTEK Spark-3.0 0.4612 0.3913 0.4025
Qwen 0.5729 0.5654 0.5394
ERNIE Bot 0.6507 0.5751 0.5773
Atom-7B 0.6385 0.3905 0.4441
Atom-1B 0.0007 0.0014 0.0009
Table 4: The Results of LLMs in Question Answering Task. "Choice" represents the score for multiple-choice questions, "Name" denotes the score for name-based question answering, "Open" corresponds to the score for open-ended question answering, and "Total" is the weighted score obtained according to the proportion of questions, representing the model’s overall effectiveness.
Model Total Choice Name Open
GPT-4 0.6850 0.9143 0.3571 0.8000
ChatGLM-4 0.7400 0.8714 0.4714 0.9000
MiniMax-ABAB-5.5 0.8250 0.8714 0.7429 0.8667
Baichuan2 0.7650 0.7857 0.6857 0.8333
XVERSE-2-65B 0.7550 0.8143 0.5857 0.8833
XVERSE-2-13B 0.6300 0.7429 0.3000 0.8833
iFLYTEK Spark-3.0 0.5400 0.7000 0.2714 0.6667
Qwen 0.7550 0.7429 0.7286 0.8000
ERNIE Bot 0.7050 0.8714 0.4000 0.8667
Atom-7B 0.4950 0.4429 0.3143 0.7667
Atom-1B 0.2500 0.4143 0.0714 0.2667
Table 5: The Results of LLMs in Text Generation Task.

### 21 Mar 2024 v3

@@ 2.1 Large Language Model
To ensure that LLMs align with human preferences and values, Reinforcement Learning from Human Feedback (RLHF) strategies are often incorporated during the training process. RLHF employs reinforcement learning algorithms, such as Proximal Policy Optimization[4], to adapt LLMs to human feedback. This method integrates human judgment into the training loop, aiming to develop LLMs that are better aligned with human values, exemplified by models like InstructGPT[5].
@@ 3 Method
In China, public security constitutes the core of police work, encompassing critical areas such as crime prevention, case investigation, intelligence analysis, and the maintenance of public order. This significant responsibility is typically borne by various specialized police forces to ensure targeted and efficient maximization. For instance, narcs play an indispensable role in drug control efforts, with their primary tasks being the identification of individuals, locations, and items related to drug activities, thereby advancing drug control initiatives. Cyber police focus on maintaining order in cyberspace, analyzing and assessing the public’s emotional responses to social events, timely identifying and dealing with online rumors, and ensuring the stability and safety of the online environment. criminal police, security officers, and traffic police play crucial roles in the investigation of criminal cases, handling civil disputes, and the control of traffic violations, respectively. Furthermore, in response to the severe challenges posed by the emerging crime of telecommunications fraud, the criminal police department has established an anti-fraud center. Its main tasks include quickly filtering potential fraudulent information, effectively issuing warnings, and blocking telecommunications fraud activities to protect the public’s financial safety and reduce economic losses. This collaborative division of labor ensures the comprehensiveness and professionalism of public security work, providing a solid foundation for maintaining social stability and citizen safety.
@@ 3.1 Tasks and Datasets
To evaluate the application performance of LLMs in the public security domain thoroughly and comprehensively, CPSDBench has constructed a four-dimensional assessment task system, encompassing key dimensions such as text classification, information extraction, question-answering, and text generation. This assessment system closely revolves around actual public security business scenarios, aggregating a diverse dataset (as shown in Table 1), aiming to conduct a more detailed and comprehensive exploration of LLMs’ abilities in parsing and generating public security-related data. Furthermore, the correlation between the evaluation content adopted by CPSDBench and the business domains of the Chinese police is graphically presented (Figure 1). Considering the high costs associated with utilizing commercial-grade large language models (e.g., ChatGPT, ChatGLM-130B, etc.), the scale of each dataset is generally controlled between tens to hundreds of samples to ensure the evaluation’s economic viability and operability.
Figure 1: The evaluative content of CPSDBench and its correspondence to the police category in China. For various police categories, the CPSDBench benchmark testing framework has designed at least five types of tasks, aiming to achieve a comprehensive and specific assessment of each category, ensuring the evaluation is both holistic and targeted.
Table 1: The evaluation dimensions and task details of CPSDBench. This benchmark framework primarily conducts an in-depth assessment of the capabilities of LLMs in the Chinese public security domain, covering four core dimensions: text classification, information extraction, question-answering, and text generation.
Evaluation Dimensions Tasks Samples Category Information
Text Classification Sentiment Classification 600 3 classes
Rumor Detection 400 2 classes
Telecommunication Fraud Detection 400 2 classes
Police Incident Classification 400 4 classes
Information Extraction Drug-Related Case NER 200 5 types of entities
Drug-Related Case RE 200 4 types of relationships
Telecommunication Fraud Case NER 200 6 types of entities
Question Answering Police Cases Reading Comprehension 200 3 types of questions
Police General Knowledge Test 200 3 types of questions
Text Generation Police Cases Summary 200 N/A
Police Cases Reports Generation 100 N/A
Police Enforcement Cases and Legal Analysis 50 N/A
@@ 3.1.1 Text Classification
  • 4)
    Police Incident Classification: Police departments routinely handle a large volume of emergency reports, and efficient classification of these reports is crucial. CPSDBench has introduced a set of real police incident data from a police command center, with theobjective of evaluating the performance of LLMs in police incident classification. The dataset involves four types of incidents: theft, telecommunication fraud, traffic cases, and intentional injury cases. Furthermore, The dataset comprises a total of 400 samples, and the distribution of the data is balanced.
@@ 3.1.2 Information Extraction
  • 1)
    Drug-Related Case Reports Information Extraction: This dataset contains report texts related to drug cases, intended for use in Named Entity Recognition (NER) and Relation Extraction tasks. For the NER task, the dataset includes five types of entities: names, dates, locations, drug types, and quantities of drugs. Notably, the data labels contain a large number of duplicate entities, making the evaluation process more stringent and complex. A prediction is considered entirely correct only when the text and quantity of an entity are accurately extracted. For the Relation Extraction task, the dataset encompasses four types of relationships: traffic_in, sell_drugs_to, possess, and provide_shelter_for.
  • 2)
    Telecommunication Fraud Case Information Extraction: This dataset is specifically designed for the identification of named entities in telecommunications fraud cases. The current version encompasses six types of entities: the reporter, the time of the incident, the location of the incident, the amount involved, social media accounts, and telephone numbers. The design of this dataset aims to evaluate the accuracy and efficiency of LLMs in processing complex textual information related to telecommunications fraud.
@@ 3.1.3 Question Answering
  • 1)
    Police Case Reading Comprehension Dataset: This dataset compiles reading comprehension questions based on actual police case files, covering both objective and subjective reading comprehension sections. For the objective part, LLMs are required to analyze the questions posed based on the provided case materials and offer answers in the form of "yes," "no," or "unknown." In the subjective reading comprehension section, LLMs need to synthesize information directly related to the question from the case materials, such as names of individuals, locations, times, and actions, to form appropriate responses. This dataset encompasses 70 objective questions, 70 subjective questions requiring high precision, and 60 subjective questions with relatively lower precision requirements.
  • 2)
    Police General Knowledge Test: The data is derived from the Chinese Police Law Enforcement Qualification Examination, aimed at assessing the LLMs legal knowledge and law enforcement capabilities related to policing, to determine their potential to assist police officers of various categories. It includes 70 true-false questions, 70 single-choice questions, and 60 multiple-choice questions.
@@ 3.1.4 Text Generation
  • 1)
    Police Case Summary Dataset: This dataset focuses on generating concise and precise summaries of police cases, aimed at presenting key information in a succinct and clear manner. Its fundamental purpose is to evaluate the model’s ability to effectively integrate and summarize essential case details.
  • 2)
    Police Cases Reports Generation: This task provides a range of basic information on public security cases, enabling LLMs to generate clear, accurate, and consistent comprehensive police reports based on this information. The core objective of this dataset is to conduct an in-depth evaluation of the quality of generation and readability of texts by LLMs in the context of producing public security reports, with the aim of enhancing the informational value of these reports. This, in turn, is expected to provide more effective support for the decision-making processes of law enforcement agencies and aid in the public’s understanding and response.
  • 3)
    Police Enforcement Cases and Legal Analysis: This dataset collects public security cases from real-world scenarios, aiming to require LLMs to conduct in-depth analyses of the provided cases and related legal provisions, and on this basis, generate solutions for problem-solving. The purpose of designing this dataset is to assess the abilities of LLMs in understanding complex law enforcement scenarios, legal logic analysis, and applying legal knowledge to solve practical problems.
@@ 3.2 Baselines and Prompt
Model Organization Size (#P) URL Access mode
GPT-3.5-turbo OpenAI $\sim$ chat.openai.com API
GPT-4 OpenAI $\sim$ chat.openai.com API
Gemini-pro Google $\sim$ gemini.google.com API
ChatGLM-4 THUDM & Knowledge Atlas 130B chatglm.cn API
MiniMax-ABAB-5.5 MiniMax $\sim$ api.minimax.chat API
Baichuan2[30] Baichuan Inc. $\sim$ www.baichuan-ai.com API
XVERSE XVERSE 13B,65B chat.xverse.cn API
iFLYTEK Spark-3.0 iFLYTEK CO.LTD. $\sim$ xinghuo.xfyun.cn API
Qwen[31] Alibaba.com Corporation 72B tongyi.aliyun.com/qianwen API
ERNIE Bot[32] Baidu, Inc. $\sim$ yiyan.baidu.com API
SenseChat SenseTime $\sim$ chat.sensetime.com API
Atom (LLaMA-2-Chinese) Llama Family 1B,7B llama.family API
@@ 3.3 Evaluation Metrics
In the objective reading comprehension component of the question-answering task, this study adopts Accuracy as the primary evaluation metric to quantify the model’s performance in identifying the correct answers. For generative question-answering and text generation tasks, the study selects Bilingual Evaluation Understudy (BLEU) and Recall-Oriented Understudy for Gisting Evaluation (ROUGE) as assessment metrics. These metrics are utilized to evaluate the model’s performance in terms of the quality and accuracy of generated text. The specific computational processes are defined in Equations ( 7) and ( 8), respectively. Equation ( 7) elaborates on the calculation method for the BLEU metric, focusing primarily on the lexical overlap between the generated text and the reference text. Equation ( 8) describes the calculation process for the ROUGE metric, emphasizing the extent to which the generated text covers and retains the content of the original text. Together, these two metrics provide a comprehensive quantitative basis for evaluating the performance of models in generative question-answering and text generation tasks.
@@
$Precision_{BERT}=\frac{1}{|C|}\sum_{c\in C}\max_{r\in R}\text{sim}(c,r)$ (9)
Furthermore, we have also selected the BERT Score[33], calculated based on the BERT-base model, as the evaluation criterion for subjective reading comprehension questions. This approach aims to measure the semantic similarity between LLMs’ predictions and the labels from a semantic perspective.The computation of BERTScore involves three primary metrics: Precision, Recall, and the F1 Score, as illustrated in equation( 9)-( 11). Here, $C$ represents the set of words in the candidate sentence, while $R$ denotes the set of words in the reference sentence. The term $sim(c,r)$ measures the similarity between the candidate word $c$ and the reference word $r$. Specifically, in subsequent experiments, we utilized the $F1_{BERT}$ as a evaluation metric for text generation tasks.
$Recall_{BERT}=\frac{1}{|R|}\sum_{r\in R}\max_{c\in C}\text{sim}(r,c)$ (10)
$F1_{BERT}=\frac{2\cdot Precision_{BERT}\cdot Recall_{BERT}}{Precision_{BERT}+Recall_{BERT}}$ (11)
@@ 4.1 Performance Evaluation
In this study, we conducted a comprehensive evaluation of advanced LLMs to explore their performance across different natural language processing tasks. The results are shown in Table 3-Table 8.Through systematic assessments in four key tasks: text classification, information extraction, question answering, and text generation, our aim was to unveil the strengths and limitations of mainstream LLMs in public security tasks.To achieve a more intuitive and comprehensive comparison of the performance of LLMs, this study also utilized radar charts to visually present the overall efficacy of the models, as shown in Figure 3. When invoking APIs to access LLMs, we ensure, as much as possible, the consistency of key parameters such as temperature, top-k, and top-p.
@@
This study further delves into the impact of language specificity on model performance. It particularly highlights that models focusing on Chinese, such as ChatGLM-4, Minimax-abab, and Baichuan2, exhibit superior performance in text generation and question-answering tasks compared to GPT-4. This finding underscores the distinct advantage of models pre-trained on rich Chinese corpora in handling tasks specific to the Chinese language. Such Chinese corpora enable LLMs to effectively learn key knowledge within the Chinese domain, such as public safety laws and law enforcement procedures, thereby proving the importance of language-specific pre-training in enhancing model performance in specific linguistic contexts.
The difficulty level of tasks also has a decisive impact on the performance of LLMs. For instance, in the task of relation extraction, models are first required to identify entities within the text and then analyze the relationships between these entities. This complexity makes such tasks significantly more challenging than text classification or entity recognition tasks, resulting in relatively lower scores for baseline models in these areas. At the same time, multiple-choice questions, compared to single-choice questions, offer a broader range of answer options, increasing the likelihood of LLMs missing or selecting too many options, thereby lowering their scores on such question types. Moreover, police enforcement cases and legal analysis tasks demand higher capabilities in text generation fluency, case logic reasoning, and mastery of public safety-related legal knowledge. The results indicate that only LLMs trained on Chinese-specific domain corpora achieve better scores in these tasks.
@@
Finally, input length is an important factor affecting the performance of LLMs in public security tasks. We observed the variation in F1 scores of LLMs when processing texts of different lengths in the task of extracting information from drug-related cases, as illustrated in Figure 1. The analysis indicates that for the majority of models, the predictive capability of LLMs tends to decline with an increase in the length of the input text.
Figure 3: A comparative analysis of the overall performance of mainstream LLMs on CPSDBench.For information extraction, the F1 score is used as the final score. The text generation task employs the BERT score as its final score. CLF refers to text classification tasks, IE denotes relation extraction tasks, QA represents question-answering tasks, and TG signifies text generation tasks.
@@
Model Sentiment Classification Rumor Detection Tel. Fraud Detection Police Incident Classification
GPT-3.5-turbo 0.6817 0.7050 0.6075 0.5300
GPT-4 0.7533 0.8975 0.8600 1.000
Gemini-pro 0.7217 0.8575 0.7150 0.9925
ChatGLM-4 0.7250 0.5750 0.8400 0.7800
MiniMax-ABAB-5.5 0.7067 0.8375 0.8050 0.9950
Baichuan2 0.6933 0.8350 0.8025 1.000
XVERSE-2-65B 0.7100 0.8175 0.4925 0.0925
XVERSE-2-13B 0.6733 0.7475 0.6425 0.9675
iFLYTEK Spark-3.0 0.6550 0.7175 0.0025 0.9875
Qwen 0.4300 0.8500 0.7125 0.8300
ERNIE Bot 0.5550 0.7575 0.6125 0.8050
SenseChat 0.7183 0.6900 0.7525 0.7375
Atom-7B 0.4683 0.7050 0.6025 0.5350
Atom-1B 0.0516 0.0600 0.4750 0.010
Table 4: The Results of LLMs in Information Extraction Tasks.
Model Drug Case NER Tel. Fraud Case NER Drug Case RE
Precision Recall F1 Score Precision Recall F1 Score Precision Recall F1 Score
GPT-3.5-turbo 0.7072 0.5952 0.6045 0.6967 0.6870 0.6845 0.1777 0.2243 0.1907
GPT-4 0.6721 0.6352 0.6295 0.7218 0.7224 0.7147 0.2533 0.3401 0.2809
Gemini-pro 0.6792 0.5715 0.5873 0.6873 0.6952 0.6830 0.1534 0.1573 0.1456
ChatGLM-4 0.5939 0.5049 0.5136 0.6347 0.6453 0.6310 0.2216 0.3548 0.2643
MiniMax-ABAB-5.5 0.6108 0.5078 0.5159 0.5805 0.5623 0.5626 0.2534 0.3610 0.2823
Baichuan2 0.6566 0.5556 0.5644 0.7042 0.6996 0.6909 0.1344 0.1896 0.1507
XVERSE-2-65B 0.6778 0.5198 0.5526 0.7149 0.6281 0.6530 0.1188 0.1642 0.1309
XVERSE-2-13B 0.5937 0.4687 0.4942 0.7136 0.6806 0.6872 0.0233 0.0292 0.0248
iFLYTEK Spark-3.0 0.4612 0.3913 0.4025 0.6053 0.5923 0.5925 0.1655 0.3210 0.2117
Qwen 0.5729 0.5654 0.5394 0.6988 0.7009 0.6905 0.1638 0.1933 0.1703
ERNIE Bot 0.6507 0.5751 0.5773 0.7050 0.7127 0.7014 0.0017 0.0017 0.0017
SenseChat 0.4211 0.3553 0.3587 0.5184 0.4644 0.4776 0.0764 0.1482 0.0973
Atom-7B 0.6385 0.3905 0.4441 0.3564 0.2698 0.2950 0.0000 0.0000 0.0000
Atom-1B 0.0007 0.0014 0.0009 0.0396 0.0285 0.0312 0.0000 0.0000 0.0000
Table 5: The Results of LLMs in Question Answering Tasks. (1)"T/F/U" and "T/F" respectively stand for True-False-Unknown questions and True-False questions. (2)"Exactitude" refers to those categories of questions that demand higher precision in the outcomes, primarily revolving around queries related to personal names. In assessing the answers to these types of questions, more stringent evaluation criteria are employed. (3)"Fuzzy" denotes other types of question and answer formats, encompassing queries related to locations, times, and methods. These queries utilize more lenient evaluation criteria (such as fuzzy matching) when evaluating responses. (4)"SCQ" and "MCQ" respectively stand for Single Choice Questions and Multiple Choice Questions. (5)"Total" is the weighted score obtained according to the proportion of questions, representing the model’s overall effectiveness.
Model Police Cases Reading Comprehension Police General Knowledge Test
Total T/F/U Exactitude Fuzzy Total T/F SCQ MCQ
GPT-3.5-turbo 0.6500 0.7000 0.4143 0.8667 0.3000 0.4429 0.3857 0.0333
GPT-4 0.6850 0.9143 0.3571 0.8000 0.4750 0.5143 0.6143 0.2667
Gemini-pro 0.8000 0.8286 0.7000 0.8833 0.3950 0.4714 0.5286 0.1500
ChatGLM-4 0.7400 0.8714 0.4714 0.9000 0.5350 0.3000 0.8429 0.4500
MiniMax-ABAB-5.5 0.8250 0.8714 0.7429 0.8667 0.5900 0.6286 0.7571 0.3500
Baichuan2 0.7650 0.7857 0.6857 0.8333 0.4750 0.6714 0.5143 0.2000
XVERSE-2-65B 0.7550 0.8143 0.5857 0.8833 0.4900 0.6143 0.5714 0.2500
XVERSE-2-13B 0.6300 0.7429 0.3000 0.8833 0.4150 0.4857 0.4857 0.2500
iFLYTEK Spark-3.0 0.5400 0.7000 0.2714 0.6667 0.5300 0.5429 0.6429 0.3833
Qwen 0.7550 0.7429 0.7286 0.8000 0.6600 0.5571 0.8571 0.5500
ERNIE Bot 0.7050 0.8714 0.4000 0.8667 0.5150 0.2714 0.7714 0.5000
SenseChat 0.5300 0.7286 0.1429 0.7500 0.3800 0.3571 0.5857 0.1667
Atom-7B 0.4950 0.4429 0.3143 0.7667 0.3100 0.4286 0.2857 0.2000
Atom-1B 0.2500 0.4143 0.0714 0.2667 0.2350 0.3857 0.2714 0.0167
Table 6: The Results of LLMs in the Police Cases Summary Task.
@@
Table 7: The Results of LLMs in the Police Cases Reports Generation Task.
GPT-3.5-turbo 0.8051 0.3103 0.2040 0.7046 0.6022 0.6257 0.5386
GPT-4 0.7539 0.2525 0.1546 0.6203 0.4882 0.4881 0.4019
Gemini-pro 0.7962 0.3298 0.2158 0.6896 0.5717 0.5917 0.4960
ChatGLM-4 0.7777 0.3113 0.2007 0.6638 0.5393 0.5461 0.4512
MiniMax-ABAB-5.5 0.7871 0.3216 0.2146 0.6789 0.5704 0.5822 0.4986
Baichuan2 0.7965 0.3415 0.2260 0.6942 0.5732 0.5932 0.4949
XVERSE-2-65B 0.7393 0.2082 0.1144 0.5605 0.4142 0.4165 0.3237
XVERSE-2-13B 0.7472 0.2368 0.1371 0.5864 0.4476 0.4516 0.3610
iFLYTEK Spark-3.0 0.7794 0.3090 0.2063 0.6560 0.5395 0.5521 0.4670
Qwen 0.7597 0.2581 0.1551 0.6102 0.4750 0.4837 0.3866
ERNIE Bot 0.7674 0.2776 0.1778 0.6312 0.5079 0.5132 0.4202
SenseChat 0.7359 0.2536 0.1593 0.5788 0.4635 0.4717 0.3901
Atom-7B 0.7253 0.2369 0.1586 0.5489 0.4690 0.4966 0.4304
Atom-1B 0.4896 0.0287 0.0183 0.1018 0.0658 0.0821 0.0582
Table 8: The Results of LLMs in the Police Enforcement Cases and Legal Analysis Task.
GPT-3.5-turbo 0.6588 0.0492 0.0200 0.3458 0.2496 0.2433 0.1909
GPT-4 0.6890 0.0973 0.0441 0.4181 0.2929 0.2954 0.2155
Gemini-pro 0.6469 0.0620 0.0335 0.3150 0.2343 0.2397 0.1908
ChatGLM-4 0.7098 0.1425 0.0752 0.4601 0.3380 0.3535 0.2710
MiniMax-ABAB-5.5 0.7275 0.1938 0.1252 0.4946 0.3985 0.3948 0.3285
Baichuan2 0.7482 0.2394 0.1665 0.5362 0.4452 0.4377 0.3759
XVERSE-2-65B 0.6998 0.1541 0.0875 0.4455 0.3267 0.3296 0.2597
XVERSE-2-13B 0.6841 0.1028 0.0530 0.4061 0.2901 0.2946 0.2216
iFLYTEK Spark-3.0 0.7016 0.1331 0.0794 0.4429 0.3309 0.3379 0.2655
Qwen 0.7529 0.2542 0.1917 0.5343 0.4453 0.4533 0.3977
ERNIE Bot 0.7626 0.2660 0.2023 0.5505 0.4565 0.4699 0.4026
SenseChat 0.6491 0.1275 0.0827 0.3329 0.2578 0.2656 0.2157
Atom-7B 0.6756 0.0823 0.0481 0.3781 0.2918 0.2871 0.2250
Atom-1B 0.5910 0.0212 0.0107 0.1941 0.1327 0.1489 0.1147